OpenStudy (anonymous):

Two cars leave town, one driving north and the other east. They are 26 miles apart when one of them is 14 miles farther from town than the other. At that time, how far were they each from the town? Distance driven by the two cars: ___miles?

OpenStudy (curry):

one of them is 6 and the other is 20

OpenStudy (anonymous):

Draw out a visual, it'll help

OpenStudy (curry):

dont forget medal

OpenStudy (anonymous):

LOL What a troll.

OpenStudy (curry):

together

OpenStudy (anonymous):

Let's do it

OpenStudy (anonymous):

And I believe you should try drawing out a triangle... with the hypotenuse and legs. That might help.

OpenStudy (anonymous):

don't worry, I always give medals for right answers, who doesn't? Let me check it out...

OpenStudy (anonymous):

it's not right.

OpenStudy (anonymous):

What isn't right?

OpenStudy (anonymous):

6, 20 is incorrect

OpenStudy (anonymous):

I know it isn't. Draw an arrow going north and one east

OpenStudy (anonymous):

Then they are 26 miles apart.Right? Make that the hypotenuse

OpenStudy (anonymous):

one car is 14 away.. so find the missing side

OpenStudy (phi):

Car A is 14 miles further from town than the other, Car B. \[ \] let the distance of Car B = x. the distance of Car A from town is 14+x use pythagoras x^{2} + (14+x)^{2} = 26^{2} which is a bit messy to solve, but doable with a calculator... \[ \] Another way, is to have memorized that a right triangle can have legs 5,12,13 or any multiple. 26 is 2*13, so you might try guessing one leg is 2*5=10, and the other leg 2*12=24. And yes, this guess works.

OpenStudy (anonymous):

Should be 21.91

OpenStudy (anonymous):

so you're saying one is 21 m and the other is 91 m? That's not right either.

OpenStudy (anonymous):

Did the question mean that when they end up driving the exact distance between them (hypontenuse) is 26?

OpenStudy (anonymous):

I think I misread the question. Oh I think it meant Leg1: x Leg2: x+14 Hypotenuse:26

OpenStudy (anonymous):

So what phi said

OpenStudy (anonymous):

alright, these wrong answers are costing me...I think you may be more experienced with high school level algebra? thanks for trying though.

OpenStudy (anonymous):

Lol then why are you asking o_o?